OpenStudy (anonymous):

7 sin^2 θ−15sin θ + 2 = 0

OpenStudy (anonymous):

7X^2 -15X +2=0 X= -(-15) + 13/ 14=2 X=-(-15)-13/14=1/7 THEN PUT X VALUE IN SINE& GET ANS.